40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 7 14:35:44 2015, diego, Update, SUS, SUS Drift Monitor 150107_SUS_DRIFTMON_screen.png
    Reply  Thu Jan 8 23:02:30 2015, diego, Update, SUS, SUS Drift Monitor 150108_SUS_DRIFTMON_screen.png
       Reply  Tue Jan 20 19:32:06 2015, Jenne, Update, SUS, SUS Drift restore scripts 
          Reply  Fri Jan 23 10:07:15 2015, Steve, Update, SUS, SUS Drift ETMX vs ETMY SUSdrift30d.png
             Reply  Mon Mar 2 18:34:12 2015, Jenne, Update, SUS, ETMX moving during single arm lock ETMX_moving.pngETMXmoving_ALSlock.pngStartsMoving_stopsItself_ranASS.png80mHz.pngIndependFromFSStemp.png
                Reply  Tue Mar 3 09:42:40 2015, Steve, Update, SUS, ETMX is drifting ETMXdrifting.pngETMXvsTemp.png
             Reply  Fri Sep 25 14:52:08 2015, Steve, Update, SUS, ETMX is not drifting ETMXstoppedDrifting.png25daysArmsT.png
                Reply  Fri Sep 25 19:06:13 2015, rana, Update, SUS, ETMX IS drifting 
Message ID: 11088     Entry time: Mon Mar 2 18:34:12 2015     In reply to: 10934     Reply to this: 11092
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMX moving during single arm lock 

This has been edited several times over the last several hours, as I try to change different parameters, to see if they affect the movement of ETMX.  So far, I don't know what is causing the motion.  If it is there, it is only present when the LSC is engaged, so I don't think it's wobbling constantly on a twisted wire.

FINAL EDIT, 9:10pm:  The arm ASC was turning itself on when the arms were locked.  Whelp, that was only 3 hours of confusion.  Blargh.

For his penance for leaving the arm ASC engaged, Q has made a set of warning lights on the LSC screen, right next to the ASS warning lights.


ETMX might be having one of those days today, which is lame. 

So far tonight, I have run the LSC offset script, set the FSS slow value to +0.2, and run the arm ASS scripts.  Nothing too crazy I think.

Sometimes when I lock the single arms, the ETMs move around like crazy.  Other times, not.  What is going on here???  The ETMs don't move at all when they are not being actuated on with the LSC.

In this screenshot you can see the end of a POX/POY lock stretch where everything was nice and good.  Then, the arms were unlocked, and they have a bit of a DC offset.  After settling from that step, they continue sitting nice and still.  Then, I relock the cavities on POX and POY a little before -4 minutes.  ETMY takes a moment to pull itself together, but then it's steady.  ETMX just wobbles around for several minutes, until I turn off the LSC enable switch (happened after the end of this plot). 

I'm not going to be able to lock like this.  Eeek! 


This is somehow related to light being in the Xarm.  This next plot was taken while the arms were held with ALS in CARM/DARM mode. 

  • Between -4 and -3, and again between -2 and -1.5, the arms were held with ALS "on resoanance".  Yarm power (green) is fine, Xarm power (purple) is not fine.  The ETMX OSEM pitch and yaw signals are bumpy at the same times.
  • Between -3 and -2 minutes, the arms were held with ALS off resonance.  Obviously you can't see anything about IR transmission, but you can see that the SUSPIT and SUSYAW signals quiet down. 
  • Between -1 and 0, I moved the Yarm ALS offset such that only it was resonating (the jags at the beginning of the resonance is me finding it, not to do with ETMY misbehaving). 
    • The ETMX SUS signals stay quiet. 
    • ETMY YAW has a DC offset when there is IR power in the arm.
  • After 0 minutes on this scale, I move the offsets such that only Xarm is "on resonance". 
    • The ETMX osem signals go bumpy again.


I closed and re-opened all 3 green shutters.  Now (at least the last 8 arm locks in the last 6 mintues) ETMX has never gone wobbly, except for a little bit right after acquisition, to deal with whatever the DC offset it. Why is this changing? 


The arms were fine for one long ~30 minute lock while I stepped out for dinner.  At some point after returned, the MC lost lock. When the arms came back, ETMX was being fussy again.  Then, it decided that it was done.

In this plot, at -1 minute I started the ASS.  Other than that, I did not touch any buttons at all, just observed.  I have no idea why at about -3 minutes the bad stuff seems to go away. 

I was curious if it had to do with the DC pointing of the optics, so I unlocked the arms, put ETMX about where it was during the long good lock stretch, then reaquired lock.  I had to undo a little of that so that it would lock on TEM00, but at the beginning of the lock stretch (starting at about -3) the pitch is about the same spot.  But, the oscillations persist.  This time it was clear that the oscillations were around 80 mHz, and they started getting bigger until they settled to an amplitude they seemed to like.


Seems pretty independent from FSS temp.  There are 3 lock stretches in the next plot (easier to see by looking at the Yarm transmission, green trace).  The first one, the FSS slow was at 0.35.  the middle one, it was around 0.05.  The last one, it was around -0.4.  Other than the different DC pointings (which I don't know if they are related), I don't see anything qualitatively different in the movement of ETMX.

Attachment 1: ETMX_moving.png  70 kB  | Hide | Hide all
ETMX_moving.png
Attachment 2: ETMXmoving_ALSlock.png  47 kB  | Hide | Hide all
ETMXmoving_ALSlock.png
Attachment 3: StartsMoving_stopsItself_ranASS.png  44 kB  Uploaded Mon Mar 2 19:52:58 2015  | Hide | Hide all
StartsMoving_stopsItself_ranASS.png
Attachment 4: 80mHz.png  48 kB  Uploaded Mon Mar 2 19:55:57 2015  | Hide | Hide all
80mHz.png
Attachment 5: IndependFromFSStemp.png  51 kB  Uploaded Mon Mar 2 20:12:58 2015  | Hide | Hide all
IndependFromFSStemp.png
ELOG V3.1.3-